hi, how do you write an equation in slope intercept form of the line that passes through the points when the points are (6.22, -3.75) and (-1.78, 0.26) ? i tried but i just keep getting it wrong when my teacher corrects it. im struggling. i forget steps my teacher says but i don't know which ones. see first i find the slope. then i put it in y=mx+b form and it never comes out right
Answer by shree840(260)

find slope m=(y2-y1)/(x2-x1)
once you find slope put in y=mx+b
yand x value you can take any point m you just found out
substitue and find b
equation is y=mx+b put values of m and b ANS